Clara earned 920 dollars during Spring break. She invested part of her money at 2 \% and the rest at 3 \%. If she received a total of 23.6 dollars in interest at the end of the year, how much was invested at 3 \%?
Answer by richwmiller(17219)

x+y=920,
0.02*x+0.03*y=23.6
x=920-y
0.02*(920-y)+0.03*y=23.6
18.4-0.02y+0.03*y=23.6
0.01*y=5.2
y=520.0 at 3%
x=920-y
x=400.0 at 2%
check
0.02*400.0+0.03*520.0=23.6
8.0+15.6=23.6
23.6=23.6
ok